OpenStudy (anonymous):

two dice are thrown prob(atleast 5 comes)=? and Prob(atmost 5 comes)=?

OpenStudy (anonymous):

2 out 12 or 1 out 6

OpenStudy (anonymous):

idk about the second one the way you are asking these are confusing

OpenStudy (anonymous):

@mvrc1996 how comes 2/12=1/6

OpenStudy (anonymous):

cause you reduce the fraction

OpenStudy (anonymous):

u didn't get my point how 2/12 comes

OpenStudy (anonymous):

total no of outcomes are 36

OpenStudy (anonymous):

i seeee hol don

OpenStudy (anonymous):

4/36 =1/9

OpenStudy (anonymous):

idk about the second one though

OpenStudy (anonymous):

okay, no pblm @mvrc1996

OpenStudy (anonymous):

but can u mention that 4 favourable outcomes

OpenStudy (anonymous):

2+3=5 4+1=5 3+2=5 1+4=5

OpenStudy (anonymous):

k

OpenStudy (zehanz):

This is the set of possible outcomes: 1 2 3 4 5 6 1 2 3 4 5 6 7 2 3 4 5 6 7 8 3 4 5 6 7 8 9 4 5 6 7 8 9 10 5 6 7 8 9 10 11 6 7 8 9 10 11 12 Each has probability (1/6)²=1/36. At most 5: there are 10 possibilities, so 10/36 = 5/18

OpenStudy (anonymous):

plse mention which are 10 favourable outcomes

OpenStudy (anonymous):

(1,0), (1,1), (1,2), (1,3), (1,4), (1,5), (1,6) (2,0), (2,1), (2,2), (2,3), (2,4), (2,5), (2,6) (3,0), (3,1), (3,2), (3,3), (3,4), (3,5), (3,6) (4,0), (4,1), (4,2), (4,3), (4,4), (4,5), (4,6) (5,0), (5,1), (5,2), (5,3), (5,4), (5,5), (5,6) (6,0), (6,1), (6,2), (6,3), (6,4), (6,5), (6,6)

OpenStudy (anonymous):

(1,0), (1,1), (1,2), (1,3) (2,0), (2,1), (2,2) (3,0), (3,1) (4,0)

OpenStudy (anonymous):

what about aleast 5 comes situation

OpenStudy (anonymous):

@ZeHanz i think u ahve given me the answer for the aleast situation

OpenStudy (zehanz):

Oops, I misread it. At least 5, so 5 or more. A lot of counting! Unless...we count the complement: max. 4. (counting) I counted 6 possibilities for max. 4 (look it up in my table) This means 5 or more will be 36-6=30 possibilities, so p=30/36=5/6

OpenStudy (anonymous):

prob(atmost ) = 30/36 and prob(atleast) =10/36

OpenStudy (anonymous):

But atmost ki liye to 15 favourable outcomes are cmg

OpenStudy (zehanz):

No, I think you made a mistake: p(sum <=5)=10/36 (10 possibilities). p(sum >=5)=30/36 (30 possibilities).

OpenStudy (anonymous):

what <= stands for

OpenStudy (zehanz):

<= smaller than or equal, so: at most >= greater than or equal, so: at least.

OpenStudy (anonymous):

but atmost condition ki liye , there 15 favourable outcomes (1,0), (1,1), (1,2), (1,3), (1,4),(2,0), (2,1), (2,2), (2,3),(3,0), (3,1), (3,2)(4,0), (4,1) and (5,0)

OpenStudy (zehanz):

0 is not on a die

OpenStudy (zehanz):

BTW, what is "ki liye"?

OpenStudy (anonymous):

it's hindi word

OpenStudy (zehanz):

Now I begrijp

OpenStudy (zehanz):

It is a dutch word :D

OpenStudy (anonymous):

i know , its meaning "understand"